Suppose 2.60 mol of an ideal gas of volume V 1 = 3.50 m 3 at T 1 = 290 K is

allowed to expand isothermally to V 2 = 7.00 m 3 at T 2 = 290 K. Determine:

a) the work done by the gas, [Ans = 4,350 J]

b) the heat added to the gas, and [Ans = take a guess]

c) the change in internal energy of the gas. [Ans = check definition of iso…]


1
Expert's answer
2022-01-18T10:20:28-0500

a) "Workdone = nRT \ln { V_1 \over V_2}"

"2.6 × 8.314 × 290( \ln{7 \over 3.5}) = 4385.62J"

b) since it is Isothermal, there is no internal energy. Then, "\Delta E_{int} = 0"

"\Delta E_{int} = Q-W"

"W= Q= 4385.62J"


c) "\Delta E_{int} = 0" (since it is isothermic)